OpenStudy (anonymous):

how to completely factor 6x^2 - 54?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Nvm that

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

oh right i see where i went wrong

OpenStudy (anonymous):

to factorise something, look for a factor that you can divide the equation by.

OpenStudy (anonymous):

Can you see anything you can divide by?

OpenStudy (anonymous):

Do you know how to break 52 into it's prime components?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

But I can divide 54 by 9 and get 6

OpenStudy (anonymous):

You know what a prime number is right? like 2,3,5,7,11, etc.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, that is already one part of the step

OpenStudy (anonymous):

Prime numbers are like atoms of numbers, you can break every number into little pieces until you can no longer divide them

OpenStudy (anonymous):

For example \[54 = 9*6 = 3*3*3*2\]

OpenStudy (anonymous):

There you have it broken down in the smallest way possible. Now if you do this for every number in your equation you can easily see which one you can factor out.

OpenStudy (anonymous):

\[\large \color{red} 3\color{red}*\color{red}2*x ^{2}-\color{red}3^{3}*\color{red}2\]

OpenStudy (anonymous):

ohh i understand

OpenStudy (anonymous):

how do factor it out?

OpenStudy (anonymous):

i broken it down as an example, can you see how to factor this?

OpenStudy (anonymous):

Do you understand what i did there?

OpenStudy (anonymous):

oh the 6 is supposed to be a 3

OpenStudy (anonymous):

i'll rewrite it

OpenStudy (anonymous):

\[\large \color{blue}3\color{blue}*\color{blue}2*\color{red}x^\color{red}2\color{green}-\color{red}3^\color{red}2*\color{blue}3\color{blue}*\color{blue}2 = \color{blue}3\color{blue}*\color{blue}2\left( \color{red}x^\color{red}2\color{green}-\color{red}3^\color{red}2 \right)\]

OpenStudy (anonymous):

sorry that took so long i was trying to make it clear with colors to show you what happens. So do you see what happened there or do you want me to explain in more detail?

OpenStudy (anonymous):

@luvsofi are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

could you explain further?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

so you see, when you calculate an equation, every time there is a + or a - or a = sign, you have to make sure that when you divide by anything, you divide by the same thing after every sign right?

OpenStudy (anonymous):

If that makes sense.

OpenStudy (anonymous):

for example 5+10=15 i can divide by 5 and it will say 1+2=3 and even though the equation itself is not the same it's still true. and also if i write it like 3*(1+2=3) than solving the brackets will give me the same equation, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the same concept is what i did with the colored equation. the green minus sign is the imaginary line. and we divide by the blue factors which are on both sides of the line and put brackets around everything else

OpenStudy (anonymous):

than put the blue factor back outside the brackets to make sure you still have the same thing if you would solve the brackets.

OpenStudy (anonymous):

sorry i'm just not understanding, i'll just ask my teacher. thank you anyways!

OpenStudy (anonymous):

you can in this case also factorize the factor inside the brackets one more time, by the way. can you see how?

OpenStudy (anonymous):

alright, no problem.